OpenStudy (anonymous):

An automobile engine uses 0.45 MJ of energy to accelerate a 1600 kg car from rest to 15 m/s. Calculate its efficiency. The formula for efficiency is output over input, but when I do that, I get 0.4. My teacher gives us the answers at the top and there's a 40, not a 0.4. I tried another excercise and I got 0.1, but there's only a 10 in the answers. My guess is that something in my process is making my answer two decimal places too small. What am I doing wrong?

OpenStudy (anonymous):

E(K)=1/2*M.V^2\[\rightarrow\] E(K)= .5*1600*(15)^2=1.8*10^5 J Ra(efficiency)=(.18/.45)*100=40%